OpenStudy (anonymous):

write each polar equation in rectangular form: r=-3sec(Theta) r=costheta+sintheta r=5/3costheta+8sin

OpenStudy (anonymous):

^theta after the last sin

myininaya (myininaya):

You need to use these equations: \[r^2=x^2+y^2 ; rcos(\theta)=x ; rsin(\theta)=y \]

OpenStudy (anonymous):

i did, but like it ended up like x^2+y^2= 25+9x^2r^2-30xr/64y^2 for the last one

myininaya (myininaya):

We want the equations just in terms of x and y instead of r and theta.

OpenStudy (anonymous):

i don't know what i did wrong

myininaya (myininaya):

so the last one is: \[r=\frac{5}{3}\cos(\theta)+8\sin(\theta) \]?

OpenStudy (anonymous):

no uh (5)/(3costheta+8sintheta)

myininaya (myininaya):

oh so you meant 5/(3costheta)+8sin(theta)) and not 5/3cos(theta)+8sin(theta)

OpenStudy (anonymous):

mhm 5 / (3cos+8sin)

myininaya (myininaya):

\[r= \frac{5}{3\cos(\theta)+8\sin(\theta)}\] Multiply both sides by that one fraction's denominator

myininaya (myininaya):

\[r(3\cos(\theta)+8\sin(\theta))=5 \] Distribute then substitute the equations I gave you

OpenStudy (anonymous):

3xr+8ry=5 but then you'll need to put r on one side cause you would need to square it. and the answer comes out all weird

myininaya (myininaya):

well rcos(theta) is x not just cos(theta same thing with the y deal rsin(theta) is y not just sin(theta)

OpenStudy (anonymous):

whoops how about the secant one

myininaya (myininaya):

put it in terms of cos by using an identity

myininaya (myininaya):

then put it in some form where you can use the equations I gave you

OpenStudy (anonymous):

sosoososososo it would be x^3+y^2x=-3 AND uh the other one x^2+y^2=x+y?

myininaya (myininaya):

for the first one? no...

myininaya (myininaya):

I haven't looked at the second one... You are making this way too hard it looks like.

myininaya (myininaya):

first one you did this as first step right?: \[r=\frac{-3}{\cos(\theta)}\]

OpenStudy (anonymous):

yep

myininaya (myininaya):

Notice this only has a cos in it... how do you get cos with the r over there so you can use the equation rcos(theta)=x

OpenStudy (anonymous):

multiply both sides with r

myininaya (myininaya):

no

myininaya (myininaya):

try again

myininaya (myininaya):

How do you undo multiplication?

myininaya (myininaya):

Or division?

OpenStudy (anonymous):

oh omg

myininaya (myininaya):

we have division by cos to undo that you ?

OpenStudy (anonymous):

x=-3

myininaya (myininaya):

yes

OpenStudy (anonymous):

thanks

myininaya (myininaya):

so for that second one...

myininaya (myininaya):

you actually got it right

myininaya (myininaya):

but how did you go about it you didn't choose a complicated way right?

OpenStudy (anonymous):

yep uh multiplied both side with r and then substituted the stuff in

myininaya (myininaya):

oh good

myininaya (myininaya):

do you have any other questions? do you think you got this better?

OpenStudy (anonymous):

uh yea i have a few more

OpenStudy (anonymous):

for standard forms of polar equation uh how would graph r=2cos3theta

myininaya (myininaya):

Plug in values of theta and see what the r output is

myininaya (myininaya):

then remember how you graph using polar coordinates r is the distance from the origin and direction is given by r to from the angle theta is the angle

OpenStudy (anonymous):

like how would you do this with a graphing calculator

myininaya (myininaya):

So for example if I plug in theta=0 \[r=2 \cos (3 \cdot 0)=2\cos(0)=2(1)=2 => \text{ graph the point } (2,0^o)\] 

myininaya (myininaya):

find 0 degrees and then go to the circle with radius 2

OpenStudy (anonymous):

cause wksht says not to use a table

myininaya (myininaya):

 then plug in more points

OpenStudy (anonymous):

oh i know how to graph but not without plugging in points

myininaya (myininaya):

I haven't looked at a calculator in ages

myininaya (myininaya):

did you try looking up how to input polar equations into whatever calculator type you have

myininaya (myininaya):

some calculators require a difference procedure in order to do certain things

myininaya (myininaya):

find mode and choose polar or pol or polargc

myininaya (myininaya):

then just type your equation in

OpenStudy (anonymous):

u what's the website

OpenStudy (anonymous):

*uh

myininaya (myininaya):

What ?